OpenStudy (izzy18970):

I need help on a question please!!??:)\ Write and solve an equation to solve each problem refer to the poster on your right if your bill for one month was $113.74,find the number of minutes you used. 9.99 a month 25cents/min.

OpenStudy (izzy18970):

please help!!??? please????????????/

OpenStudy (solomonzelman):

I don't get the question, do you need minutes per 9.99 months of water used?

OpenStudy (izzy18970):

I think so

OpenStudy (izzy18970):

25p + 9.99 = 113.74 ??

OpenStudy (solomonzelman):

No, hold on.

OpenStudy (izzy18970):

is this the correct equation?

OpenStudy (izzy18970):

.25p + 9.99 = 113.74

OpenStudy (solomonzelman):

25c/per minute say 0.25 dollars per minute, to convert dollars into minutes 0.25d + 9.99 = 113.74 (in dollars not cents) $113.74 * (1/4) or 113.74 / 4 =8.35

OpenStudy (izzy18970):

i got p=415

OpenStudy (izzy18970):

8.35 doesnt = 113.74